Let us consider a quadratic polynomial f(x) such that equation f(x)=8x-16 has exactly one root, and equation f(x)=2x-4 has exactly one root. Find the maximum value of a discriminant of f(x)
1
Expert's answer
2021-01-07T14:30:14-0500

"f(x) = ax^2 + bx + c"

"1.\ ax^2 + bx + c = 8x - 16"

"2. \ ax^2 + bx + c = 2x - 4"


"1. \ ax^2 + x(b -8) + c + 16 = 0"

"2.\ ax^2 + x(b - 2) + c + 4= 0"


"x_1 = x_2 \leftrightarrow D = 0;"

"D = b^2 -4ac;"

"\therefore"

"1. \ (b - 8)^2 - 4a(c +16) = 0"

"2. \ (b - 2)^2 - 4a(c + 4) = 0"


"1. \ b^2 -16b + 64 - 4ac - 64a = 0"

"2. \ b^2 -4b + 4 -4ac - 16a = 0"


Subtracting 2. from 1. yields:

"-12b + 60 -48a = 0"

"b - 5 + 4a = 0"

"b = 5 -4a \ (*)"

From this and 2. :

"(5 - 4a -2)^2 -4ac -16a = 0"

"(3 - 4a)^2 - 4ac -16a = 0"

"9 -24a +16a^2 - 4ac - 16a = 0"

"16a^2 - 40a + 9 - 4ac = 0"

"-16a^2 + 40a - 9 = -4ac"

From this and "(*)" :

"D_{f(x)} = b^2 - 4ac ="

"= (5 - 4a)^2 - 16a^2 + 40a - 9 ="

"= 25 - 40a + 16a^2 - 16a^2 + 40a - 9 = 16"


Answer: "D_{max} = 16"